3. In the hydraulic lift arrangement shown at right, the two masses sit at rest on P P. massless cylinder heads above a reservoir of incompressible hydraulic fluid. The chambers above the masses are sealed, with gas pressures P, and P2 Evaluate the following statement (T/F/N). Explain your reasoning. m2 If d2 3dh 2h,, and m,-8m, then P, must be greater than P, h2
Answer #1

let d is the density of fluid. Fluid equation at P3 and corresponding point of same pressure in first tube.

P1+dgh1+m1*4/3.14d12=P2+dgh2+m2*4/3.14d22

P2-P1=dgh2+(1/9)m1*4/3.14d12

right hand side of the equation is always positive .

hence P2-P1 is positive

or P2>P1